OpenStudy (anonymous):

How can i find the product of the complex number and its conjugate. 1 + 2i

OpenStudy (anonymous):

sorry... i mean 1+2*i??

OpenStudy (anonymous):

The question says "1+2i"

OpenStudy (anonymous):

mmmmm... lets see....

ganeshie8 (ganeshie8):

\((a+bi)(a-bi) = a^2 + b^2\)

ganeshie8 (ganeshie8):

conjugate of \(1+2i\) is \(1-2i\)

ganeshie8 (ganeshie8):

product wud be : \((1+2i)(1-2i) = 1^2 + 2^2 = 1 + 4 = 5\)

OpenStudy (anonymous):

oohhh i get it. but i dont know where she went??? well, thanks!!

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

Still here!

OpenStudy (anonymous):

Thanks!

ganeshie8 (ganeshie8):

thank ninjastrike :)

OpenStudy (anonymous):

your welcome